### v2

# Euclid Quick Data Release (Q1)

11institutetext: INAF-Osservatorio Astronomico di Trieste, Via G. B. Tiepolo 11, 34143 Trieste, Italy 22institutetext: Universitäts-Sternwarte München, Fakultät für Physik, Ludwig-Maximilians-Universität München, Scheinerstrasse 1, 81679 München, Germany 33institutetext: Université Paris-Saclay, CNRS, Institut d’astrophysique spatiale, 91405, Orsay, France 44institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstr. 1, 85748 Garching, Germany 55institutetext: INAF-Osservatorio Astronomico di Roma, Via Frascati 33, 00078 Monteporzio Catone, Italy 66institutetext: Institut de Recherche en Astrophysique et Planétologie (IRAP), Université de Toulouse, CNRS, UPS, CNES, 14 Av. Edouard Belin, 31400 Toulouse, France 77institutetext: Observatorio Nacional, Rua General Jose Cristino, 77-Bairro Imperial de Sao Cristovao, Rio de Janeiro, 20921-400, Brazil 88institutetext: Instituto de Astrofísica de Canarias, Vía Láctea, 38205 La Laguna, Tenerife, Spain 99institutetext: Instituto de Astrofísica de Canarias (IAC); Departamento de Astrofísica, Universidad de La Laguna (ULL), 38200, La Laguna, Tenerife, Spain 1010institutetext: Université PSL, Observatoire de Paris, Sorbonne Université, CNRS, LERMA, 75014, Paris, France $1$institutetext: Université Paris-Cité, 5 Rue Thomas Mann, 75013, Paris, France 1212institutetext: Dipartimento di Fisica - Sezione di Astronomia, Università di Trieste, Via Tiepolo 11, 34131 Trieste, Italy 1313institutetext: David A. Dunlap Department of Astronomy & Astrophysics, University of Toronto, 50 St George Street, Toronto, Ontario M5S 3H4, Canada 1414institutetext: Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, University of Manchester, Oxford Road, Manchester M13 9PL, UK 1515institutetext: Université Paris Cité, CNRS, Astroparticule et Cosmologie, 75013 Paris, France 1616institutetext: Department of Astronomy, University of Geneva, ch. d’Ecogia 16, 1290 Versoix, Switzerland 1717institutetext: ESAC/ESA, Camino Bajo del Castillo, s/n., Urb. Villafranca del Castillo, 28692 Villanueva de la Cañada, Madrid, Spain 1818institutetext: School of Mathematics and Physics, University of Surrey, Guildford, Surrey, GU2 7XH, UK 1919institutetext: INAF-Osservatorio Astronomico di Brera, Via Brera 28, 20122 Milano, Italy 2020institutetext: INAF-Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Piero Gobetti 93/3, 40129 Bologna, Italy 2121institutetext: Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France $2$institutetext: IFPU, Institute for Fundamental Physics of the Universe, via Beirut 2, 34151 Trieste, Italy 2323institutetext: INFN, Sezione di Trieste, Via Valerio 2, 34127 Trieste TS, Italy 2424institutetext: SISSA, International School for Advanced Studies, Via Bonomea 265, 34136 Trieste TS, Italy 2525institutetext: Dipartimento di Fisica e Astronomia, Università di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 2626institutetext: INFN-Sezione di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 2727institutetext: INAF-Osservatorio Astronomico di Padova, Via dell’Osservatorio 5, 35122 Padova, Italy 2828institutetext: Centre National d’Etudes Spatiales – Centre spatial de Toulouse, 18 avenue Edouard Belin, 31401 Toulouse Cedex 9, France 2929institutetext: Kapteyn Astronomical Institute, University of Groningen, PO Box 800, 9700 AV Groningen, The Netherlands 3030institutetext: ATG Europe BV, Huygensstraat 34, 2201 DK Noordwijk, The Netherlands 3131institutetext: Space Science Data Center, Italian Space Agency, via del Politecnico snc, 00133 Roma, Italy 3232institutetext: INAF-Osservatorio Astrofisico di Torino, Via Osservatorio 20, 10025 Pino Torinese (TO), Italy $3$institutetext: Dipartimento di Fisica, Università di Genova, Via Dodecaneso 33, 16146, Genova, Italy 3434institutetext: INFN-Sezione di Genova, Via Dodecaneso 33, 16146, Genova, Italy 3535institutetext: Department of Physics ”E. Pancini”, University Federico II, Via Cinthia 6, 80126, Napoli, Italy 3636institutetext: INAF-Osservatorio Astronomico di Capodimonte, Via Moiariello 16, 80131 Napoli, Italy 3737institutetext: Instituto de Astrofísica e Ciências do Espaço, Universidade do Porto, CAUP, Rua das Estrelas, PT4150-762 Porto, Portugal 3838institutetext: Faculdade de Ciências da Universidade do Porto, Rua do Campo de Alegre, 4150-007 Porto, Portugal 3939institutetext: Dipartimento di Fisica, Università degli Studi di Torino, Via P. Giuria 1, 10125 Torino, Italy 4040institutetext: INFN-Sezione di Torino, Via P. Giuria 1, 10125 Torino, Italy 4141institutetext: European Space Agency/ESTEC, Keplerlaan 1, 2201 AZ Noordwijk, The Netherlands 4242institutetext: Institute Lorentz, Leiden University, Niels Bohrweg 2, 2333 CA Leiden, The Netherlands 4343institutetext: Leiden Observatory, Leiden University, Einsteinweg 55, 2333 CC Leiden, The Netherlands $4$institutetext: INAF-IASF Milano, Via Alfonso Corti 12, 20133 Milano, Italy 4545institutetext: Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Avenida Complutense 40, 28040 Madrid, Spain 4646institutetext: Port d’Informació Científica, Campus UAB, C. Albareda s/n, 08193 Bellaterra (Barcelona), Spain 4747institutetext: Institute for Theoretical Particle Physics and Cosmology (TTK), RWTH Aachen University, 52056 Aachen, Germany 4848institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, 08193 Barcelona, Spain 4949institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), Edifici RDIT, Campus UPC, 08860 Castelldefels, Barcelona, Spain 5050institutetext: INFN section of Naples, Via Cinthia 6, 80126, Napoli, Italy 5151institutetext: Institute for Astronomy, University of Hawaii, 2680 Woodlawn Drive, Honolulu, HI 96822, USA 5252institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, Viale Berti Pichat 6/2, 40127 Bologna, Italy 5353institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 5454institutetext: European Space Agency/ESRIN, Largo Galileo Galilei 1, 00044 Frascati, Roma, Italy $5$institutetext: Université Claude Bernard Lyon 1, CNRS/IN2P3, IP2I Lyon, UMR 5822, Villeurbanne, F-69100, France 5656institutetext: Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (IEEC-UB), Martí i Franquès 1, 08028 Barcelona, Spain 5757institutetext: Institució Catalana de Recerca i Estudis Avançats (ICREA), Passeig de Lluís Companys 23, 08010 Barcelona, Spain 5858institutetext: UCB Lyon 1, CNRS/IN2P3, IUF, IP2I Lyon, 4 rue Enrico Fermi, 69622 Villeurbanne, France 5959institutetext: Mullard Space Science Laboratory, University College London, Holmbury St Mary, Dorking, Surrey RH5 6NT, UK 6060institutetext: Canada-France-Hawaii Telescope, 65-1238 Mamalahoa Hwy, Kamuela, HI 96743, USA 6161institutetext: Aix-Marseille Université, CNRS, CNES, LAM, Marseille, France 6262institutetext: Departamento de Física, Faculdade de Ciências, Universidade de Lisboa, Edifício C8, Campo Grande, PT1749-016 Lisboa, Portugal 6363institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Campo Grande, 1749-016 Lisboa, Portugal 6464institutetext: INAF-Istituto di Astrofisica e Planetologia Spaziali, via del Fosso del Cavaliere, 100, 00100 Roma, Italy 6565institutetext: INFN-Padova, Via Marzolo 8, 35131 Padova, Italy $6$institutetext: Aix-Marseille Université, CNRS/IN2P3, CPPM, Marseille, France 6767institutetext: INFN-Bologna, Via Irnerio 46, 40126 Bologna, Italy 6868institutetext: School of Physics, HH Wills Physics Laboratory, University of Bristol, Tyndall Avenue, Bristol, BS8 1TL, UK 6969institutetext: NRC Herzberg, 5071 West Saanich Rd, Victoria, BC V9E 2E7, Canada 7070institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O. Box 1029 Blindern, 0315 Oslo, Norway 7171institutetext: Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA, 91109, USA 7272institutetext: Department of Physics, Lancaster University, Lancaster, LA1 4YB, UK 7373institutetext: Felix Hormuth Engineering, Goethestr. 17, 69181 Leimen, Germany 7474institutetext: Technical University of Denmark, Elektrovej 327, 2800 Kgs. Lyngby, Denmark 7575institutetext: Cosmic Dawn Center (DAWN), Denmark 7676institutetext: Max-Planck-Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany $7$institutetext: NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 7878institutetext: Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 7979institutetext: Department of Physics and Helsinki Institute of Physics, Gustaf Hällströmin katu 2, 00014 University of Helsinki, Finland 8080institutetext: Université de Genève, Département de Physique Théorique and Centre for Astroparticle Physics, 24 quai Ernest-Ansermet, CH-1211 Genève 4, Switzerland 8181institutetext: Department of Physics, P.O. Box 64, 00014 University of Helsinki, Finland 8282institutetext: Helsinki Institute of Physics, Gustaf Hällströmin katu 2, University of Helsinki, Helsinki, Finland 8383institutetext: Centre de Calcul de l’IN2P3/CNRS, 21 avenue Pierre de Coubertin 69627 Villeurbanne Cedex, France 8484institutetext: Laboratoire d’etude de l’Univers et des phenomenes eXtremes, Observatoire de Paris, Université PSL, Sorbonne Université, CNRS, 92190 Meudon, France 8585institutetext: SKA Observatory, Jodrell Bank, Lower Withington, Macclesfield, Cheshire SK11 9FT, UK 8686institutetext: Dipartimento di Fisica ”Aldo Pontremoli”, Università degli Studi di Milano, Via Celoria 16, 20133 Milano, Italy 8787institutetext: INFN-Sezione di Milano, Via Celoria 16, 20133 Milano, Italy $8$institutetext: University of Applied Sciences and Arts of Northwestern Switzerland, School of Engineering, 5210 Windisch, Switzerland 8989institutetext: Universität Bonn, Argelander-Institut für Astronomie, Auf dem Hügel 71, 53121 Bonn, Germany 9090institutetext: INFN-Sezione di Roma, Piazzale Aldo Moro, 2 - c/o Dipartimento di Fisica, Edificio G. Marconi, 00185 Roma, Italy 9191institutetext: Dipartimento di Fisica e Astronomia ”Augusto Righi” - Alma Mater Studiorum Università di Bologna, via Piero Gobetti 93/2, 40129 Bologna, Italy 9292institutetext: Department of Physics, Institute for Computational Cosmology, Durham University, South Road, Durham, DH1 3LE, UK 9393institutetext: Université Côte d’Azur, Observatoire de la Côte d’Azur, CNRS, Laboratoire Lagrange, Bd de l’Observatoire, CS 34229, 06304 Nice cedex 4, France 9494institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS, and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 9595institutetext: CNRS-UCB International Research Laboratory, Centre Pierre Binetruy, IRL2007, CPB-IN2P3, Berkeley, USA 9696institutetext: Institut d’Astrophysique de Paris, 98bis Boulevard Arago, 75014, Paris, France 9797institutetext: Institute of Physics, Laboratory of Astrophysics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 9898institutetext: Aurora Technology for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain $9$institutetext: Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona), Spain 100100institutetext: School of Mathematics, Statistics and Physics, Newcastle University, Herschel Building, Newcastle-upon-Tyne, NE1 7RU, UK 101101institutetext: DARK, Niels Bohr Institute, University of Copenhagen, Jagtvej 155, 2200 Copenhagen, Denmark 102102institutetext: Waterloo Centre for Astrophysics, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada 103103institutetext: Department of Physics and Astronomy, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada 104104institutetext: Perimeter Institute for Theoretical Physics, Waterloo, Ontario N2L 2Y5, Canada 105105institutetext: Institute of Space Science, Str. Atomistilor, nr. 409 Măgurele, Ilfov, 077125, Romania 106106institutetext: Dipartimento di Fisica e Astronomia ”G. Galilei”, Università di Padova, Via Marzolo 8, 35131 Padova, Italy 107107institutetext: Institut für Theoretische Physik, University of Heidelberg, Philosophenweg 16, 69120 Heidelberg, Germany 108108institutetext: Université St Joseph; Faculty of Sciences, Beirut, Lebanon 109109institutetext: Departamento de Física, FCFM, Universidad de Chile, Blanco Encalada 2008, Santiago, Chile 110110institutetext: Satlantis, University Science Park, Sede Bld 48940, Leioa-Bilbao, Spain 111111institutetext: Department of Physics, Royal Holloway, University of London, TW20 0EX, UK 112112institutetext: Infrared Processing and Analysis Center, California Institute of Technology, Pasadena, CA 91125, USA 113113institutetext: Instituto de Astrofísica e Ciências do Espaço, Faculdade de Ciências, Universidade de Lisboa, Tapada da Ajuda, 1349-018 Lisboa, Portugal 114114institutetext: Cosmic Dawn Center (DAWN) 115115institutetext: Niels Bohr Institute, University of Copenhagen, Jagtvej 128, 2200 Copenhagen, Denmark 116116institutetext: Universidad Politécnica de Cartagena, Departamento de Electrónica y Tecnología de Computadoras, Plaza del Hospital 1, 30202 Cartagena, Spain 117117institutetext: Centre for Information Technology, University of Groningen, P.O. Box 11044, 9700 CA Groningen, The Netherlands 118118institutetext: Dipartimento di Fisica e Scienze della Terra, Università degli Studi di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 119119institutetext: Istituto Nazionale di Fisica Nucleare, Sezione di Ferrara, Via Giuseppe Saragat 1, 44122 Ferrara, Italy 120120institutetext: INAF, Istituto di Radioastronomia, Via Piero Gobetti 101, 40129 Bologna, Italy 121121institutetext: Department of Physics, Oxford University, Keble Road, Oxford OX1 3RH, UK 122122institutetext: Ernst-Reuter-Str. 4e, 31224 Peine, Germany 123123institutetext: INAF - Osservatorio Astronomico di Brera, via Emilio Bianchi 46, 23807 Merate, Italy 124124institutetext: INAF-Osservatorio Astronomico di Brera, Via Brera 28, 20122 Milano, Italy, and INFN-Sezione di Genova, Via Dodecaneso 33, 16146, Genova, Italy 125125institutetext: ICL, Junia, Université Catholique de Lille, LITL, 59000 Lille, France 126126institutetext: ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing, Via Magnanelli 2, Bologna, Italy 127127institutetext: Instituto de Física Teórica UAM-CSIC, Campus de Cantoblanco, 28049 Madrid, Spain 128128institutetext: CERCA/ISO, Department of Physics, Case Western Reserve University, 10900 Euclid Avenue, Cleveland, OH 44106, USA 129129institutetext: Technical University of Munich, TUM School of Natural Sciences, Physics Department, James-Franck-Str. 1, 85748 Garching, Germany 130130institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild-Str. 1, 85748 Garching, Germany 131131institutetext: Laboratoire Univers et Théorie, Observatoire de Paris, Université PSL, Université Paris Cité, CNRS, 92190 Meudon, France 132132institutetext: Departamento de Física Fundamental. Universidad de Salamanca. Plaza de la Merced s/n. 37008 Salamanca, Spain 133133institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, 67000 Strasbourg, France 134134institutetext: Center for Data-Driven Discovery, Kavli IPMU (WPI), UTIAS, The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 135135institutetext: Ludwig-Maximilians-University, Schellingstrasse 4, 80799 Munich, Germany 136136institutetext: Max-Planck-Institut für Physik, Boltzmannstr. 8, 85748 Garching, Germany 137137institutetext: California Institute of Technology, 1200 E California Blvd, Pasadena, CA 91125, USA 138138institutetext: University of California, Los Angeles, CA 90095-1562, USA 139139institutetext: Department of Physics & Astronomy, University of California Irvine, Irvine CA 92697, USA 140140institutetext: Department of Mathematics and Physics E. De Giorgi, University of Salento, Via per Arnesano, CP-I93, 73100, Lecce, Italy 141141institutetext: INFN, Sezione di Lecce, Via per Arnesano, CP-193, 73100, Lecce, Italy 142142institutetext: INAF-Sezione di Lecce, c/o Dipartimento Matematica e Fisica, Via per Arnesano, 73100, Lecce, Italy 143143institutetext: Departamento Física Aplicada, Universidad Politécnica de Cartagena, Campus Muralla del Mar, 30202 Cartagena, Murcia, Spain 144144institutetext: Instituto de Física de Cantabria, Edificio Juan Jordá, Avenida de los Castros, 39005 Santander, Spain 145145institutetext: Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth PO1 3FX, UK 146146institutetext: Department of Computer Science, Aalto University, PO Box 15400, Espoo, FI-00 076, Finland 147147institutetext: Instituto de Astrofísica de Canarias, c/ Via Lactea s/n, La Laguna 38200, Spain. Departamento de Astrofísica de la Universidad de La Laguna, Avda. Francisco Sanchez, La Laguna, 38200, Spain 148148institutetext: Universidad de La Laguna, Departamento de Astrofísica, 38206 La Laguna, Tenerife, Spain 149149institutetext: Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute (AIRUB), German Centre for Cosmological Lensing (GCCL), 44780 Bochum, Germany 150150institutetext: Department of Physics and Astronomy, Vesilinnantie 5, 20014 University of Turku, Finland 151151institutetext: Serco for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 152152institutetext: ARC Centre of Excellence for Dark Matter Particle Physics, Melbourne, Australia 153153institutetext: Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Hawthorn, Victoria 3122, Australia 154154institutetext: Department of Physics and Astronomy, University of the Western Cape, Bellville, Cape Town, 7535, South Africa 155155institutetext: DAMTP, Centre for Mathematical Sciences, Wilberforce Road, Cambridge CB3 0WA, UK 156156institutetext: Kavli Institute for Cosmology Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 157157institutetext: Department of Astrophysics, University of Zurich, Winterthurerstrasse 190, 8057 Zurich, Switzerland 158158institutetext: IRFU, CEA, Université Paris-Saclay 91191 Gif-sur-Yvette Cedex, France 159159institutetext: Oskar Klein Centre for Cosmoparticle Physics, Department of Physics, Stockholm University, Stockholm, SE-106 91, Sweden 160160institutetext: Astrophysics Group, Blackett Laboratory, Imperial College London, London SW7 2AZ, UK 161161institutetext: Univ. Grenoble Alpes, CNRS, Grenoble INP, LPSC-IN2P3, 53, Avenue des Martyrs, 38000, Grenoble, France 162162institutetext: INAF-Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, 50125, Firenze, Italy 163163institutetext: Dipartimento di Fisica, Sapienza Università di Roma, Piazzale Aldo Moro 2, 00185 Roma, Italy 164164institutetext: Centro de Astrofísica da Universidade do Porto, Rua das Estrelas, 4150-762 Porto, Portugal 165165institutetext: HE Space for European Space Agency (ESA), Camino bajo del Castillo, s/n, Urbanizacion Villafranca del Castillo, Villanueva de la Cañada, 28692 Madrid, Spain 166166institutetext: Department of Astrophysical Sciences, Peyton Hall, Princeton University, Princeton, NJ 08544, USA 167167institutetext: Theoretical astrophysics, Department of Physics and Astronomy, Uppsala University, Box 515, 751 20 Uppsala, Sweden 168168institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 169169institutetext: Mathematical Institute, University of Leiden, Einsteinweg 55, 2333 CA Leiden, The Netherlands 170170institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 171171institutetext: Space physics and astronomy research unit, University of Oulu, Pentti Kaiteran katu 1, FI-90014 Oulu, Finland 172172institutetext: Center for Computational Astrophysics, Flatiron Institute, 162 5th Avenue, 10010, New York, NY, USA 173173institutetext: Department of Astronomy, University of Massachusetts, Amherst, MA 01003, USA 174174institutetext: Univ. Lille, CNRS, Centrale Lille, UMR 9189 CRIStAL, 59000 Lille, France 175175institutetext: Department of Physics and Astronomy, University of British Columbia, Vancouver, BC V6T 1Z1, Canada
Euclid Quick Data Release (Q1)
From images to multiwavelength catalogues: the Euclid MERge Processing Functionthanks: We wish to remember our colleagues Mario Nonino and Stefano Pilo, who played an important role in OU-MER, and, more importantly, in our lives.
Euclid Collaboration: E. Romelli erik.romelli@inaf.it Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Kümmel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Dole Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Gracia-Carpio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Merlin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Galeotta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Fang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Castellano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Caro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Soubrie Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Maurin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Cabanac Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Dimauro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Huertas-Company Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. D. Lepinzan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Vassallo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Walmsley Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. A. Zinchenko Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Boucaud Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Calabro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Roscani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Tramacere Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Douspis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Fontana Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Aghanim Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Altieri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Amara Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Andreon Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Auricchio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Aussel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Baccigalupi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Baldi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Balestra Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Bardelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Basset Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Battaglia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. N. Belikov Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Biviano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Bonchi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Bonino Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Branchini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Brescia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Brinchmann Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Camera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Cañas-Herrera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Capobianco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Carbone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Carretero Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Casas Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. J. Castander Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Castignani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Cavuoti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. C. Chambers Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Cimatti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Colodro-Conde Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Congedo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. J. Conselice Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Conversi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Copin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Courbin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. M. Courtois Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Cropper Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J.-G. Cuby Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Da Silva Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. da Silva Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Degaudenzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. de la Torre Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. De Lucia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. M. Di Giorgio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Dubath Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. A. J. Duncan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    X. Dupac Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Dusini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Escoffier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Fabricius Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Farina Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Farinelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Faustini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Ferriol Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Finelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Fotopoulou Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Fourmanoit Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Frailis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Franceschi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. George Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. Gillard Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Gillis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Giocoli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. R. Granett Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Grazian Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Grupp Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Gwyn Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. V. H. Haugan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Hoar Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Hoekstra Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. Holmes Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. M. Hook Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Hormuth Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Hornstrup Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Jahnke Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Jhabvala Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Joachimi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Keihänen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Kermiche Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Kiessling Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Kilbinger Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Kubik Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Kuijken Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Kunz Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Kurki-Suonio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Laureijs Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Q. Le Boulc’h Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. M. C. Le Brun Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Le Mignant Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Liebing Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Ligori Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. B. Lilje Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Lindholm Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Lloro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Mainetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Maino Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Maiorano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Mansutti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Marcin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Marggraf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Martinelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Martinet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Marulli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Massey Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Maurogordato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. J. McCracken Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Medinaceli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Mei Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Melchior Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Mellier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Meneghetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Meylan Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Mora Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Moresco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Moscardini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Nakajima Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Neissner Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. C. Nichol Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S.-M. Niemi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. W. Nightingale Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Padilla Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Paltani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Pasian Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Pedersen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. J. Percival Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Pettorino Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Pires Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Polenta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Poncet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. A. Popa Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Pozzetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. D. Racca Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Raison Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Renzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Rhodes Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Riccio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Roncarelli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Rossetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Saglia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Z. Sakr Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. G. Sánchez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Sapone Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Sartoris Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. A. Schewtschenko Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Schirmer Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Schneider Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Scodeggio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Secroun Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sefusatti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Seidel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Seiffert Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Serrano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Simon Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Sirignano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Sirri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Spurio Mancini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Stanco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Steinwagner Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Tallada-Crespí Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Tavagnacco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. N. Taylor Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. I. Teplitz Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Tereno Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Tessore Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Toft Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Toledo-Moreo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Torradeflot Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Tsyganov Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Tutusaus Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. A. Valentijn Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Valenziano Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Valiviita Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Verdoes Kleijn Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Veropalumbo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Wang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Weller Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Zamorani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. M. Zerbi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Zucca Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Allevato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Ballardini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Bolzonella Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Bozzo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Burigana Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Cappi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Casenove Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Di Ferdinando Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. A. Escartin Vigo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Gabarra Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    W. G. Hartley Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Israel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Martín-Fleitas Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Matthew Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. Mauri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. B. Metcalf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Pezzotta Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Pöntinen Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Porciani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. Risso Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Scottez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Sereno Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Tenti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Viel Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Wiesmann Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Akrami Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Alvi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    I. T. Andika Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Anselmi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Archidiacono Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Atrio-Barandela Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Bergamini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Bertacca Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Bethermin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Blanchard Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Blot Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Böhringer Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Borgani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. L. Brown Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Bruton Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    B. Camacho Quevedo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. S. Carvalho Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Castro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Chary Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Cogato Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Conseil Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. R. Cooray Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    O. Cucciati Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Davini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. De Paolis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Desprez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Díaz-Sánchez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. J. Diaz Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Di Domizio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. M. Diego Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P.-A. Duc Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Enia Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. M. N. Ferguson Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Finoguenov Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Franco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Ganga Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. García-Bellido Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. Gasparetto Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Gavazzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Gaztanaga Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Giacomini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Gianotti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Gozaliasl Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Gregorio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Guidi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. M. Gutierrez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Hall Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Hernández-Monteagudo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    H. Hildebrandt Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Hjorth Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. J. E. Kajava Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    Y. Kang Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Kansal Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Karagiannis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Kiiveri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. C. Kirkpatrick Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Kruk Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    V. Le Brun Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Le Graet Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Legrand Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Lembo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Lepori Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. F. Lesci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Lesgourgues Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Leuzzi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    T. I. Liaudat Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Loureiro Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. Macias-Perez Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Maggio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Magliocchetti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. A. Magnier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Mancini Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Mannucci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Maoli Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. J. A. P. Martins Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Miluzio Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Monaco Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Moretti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Morgante Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Nadathur Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Naidoo Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Navarro-Alsina Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Nesseris Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    F. Passalacqua Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Paterson Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. Patrizii Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Pisani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Potter Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Quai Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Radovich Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Reimberg Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P.-F. Rocci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Rodighiero Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Sacquegna Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Sahlén Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. B. Sanders Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sarpa Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Scarlata Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Schneider Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Schultheis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Sciotti Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    E. Sellentin Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    L. C. Smith Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    K. Tanidis Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Testera Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    R. Teyssier Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    S. Tosi Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Troja Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    M. Tucci Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    C. Valieri Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    A. Venhola Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Vergani Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    G. Verza Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    P. Vielzeuf Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    N. A. Walton Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. R. Weaver Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    J. G. Sorce Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)    D. Scott Euclid Quick Data Release (Q1)Euclid Quick Data Release (Q1)
[…]
Key Words.:
Galaxies:photometry, Galaxies:morphology, Catalogues, Methods:data analysis
1 Introduction
The European Space Agency’s Euclid mission (Euclid Collaboration: Mellier et al. 2024) is dedicated to studying the dark Universe through a photometric and spectroscopic survey of the extragalactic sky at visible and near-infrared wavelengths. Successfully launched from Cape Canaveral on July 1 2023, the Euclid satellite aims to gather data for the Euclid Wide Survey (EWS) and the Euclid Deep Survey (EDS), both described in details in Euclid Collaboration: Scaramella et al. (2022), by means of two instruments, the Visible Camera (VIS, Euclid Collaboration: Cropper et al. 2024) and the Near-Infrared Spectrometer and Photometer (NISP, Euclid Collaboration: Jahnke et al. 2024).
To efficiently manage and process the immense volume of data generated by Euclid, the Euclid Consortium (EC) has developed a suite of software Processing Functions (PFs) forming the Euclid Science Ground Segment (SGS). Organizational Units (OUs) are responsible for developing, maintaining and validating the PFs. A detailed description on how the SGS processing is organized and how the OUs interact with each other is found in Euclid Collaboration: Aussel et al. (2025).
The MER PF realises the merging of photometric information from VIS and NISP, together with information from external photometric surveys: DES (Abbott et al. 2021), CFIS (Ibata et al. 2017), Pan-STARRS (Magnier et al. 2020), WHIGS and WHISHES (see Aihara et al. 2017), Rubin-LSST (Ivezić et al. 2019). The main objective of the MER PF is to provide source catalogues containing object information relevant for Euclid’s science goals.
The primary requirement for which the MER PF has been designed is the production of a image mosaics, and a photometric catalogue of sources. To provide this, it starts from VIS and NISP calibrated frames, processed by OU-VIS (Euclid Collaboration: McCracken et al. 2025) and OU-NIR (Euclid Collaboration: Polenta et al. 2025) respectively, and external surveys’ stacked frames, processed by OU-EXT (Euclid Collaboration: Aussel et al. 2025), together with related ancillary data and PSF models. Separate object detections are performed on both the photometric VIS and NIR (stacking the three NISP bands $J_{\scriptscriptstyle\rm E}$, $H_{\scriptscriptstyle\rm E}$, and $Y_{\scriptscriptstyle\rm E}$) mosaics. A multi-wavelength flux determination is performed on detected objects with suitable photometric techniques. The final product of the MER processing function is a catalogue of sources with unique identification number, multi-wavelength photometric measurements, and ancillary information (morphological estimates, star/galaxy separation, etc.).
The MER PF also provides the required input, as additional information integrated in the final Euclid catalogue, for photometric redshift computation and for spectra extraction for which respectively OU-PHZ (Euclid Collaboration: Tucci et al. 2025) and OU-SIR (Euclid Collaboration: Copin et al. 2025) are in charge as customer OUs of OU-MER. The MER PF catalogues represent also a primary input for shear estimation, which is performed by OU-SHE. The MER products are among the core products delivered in the first Euclid data release, Q1 (Euclid Quick Release Q1 2025; Euclid Collaboration: Aussel et al. 2025), and are used in many science papers.
2 The Euclid MER software pipeline overview
The MER PF is designed around the concept of pipeline as a chain of Processing Elements (PEs). Each PE represents an atomic block that cannot be split further from a functional point of view. The tasks related to each PE are implemented by one or more software components. Source codes are in the official Euclid GitLab repository and are available throughout the EC via the COmmon DEvelopment ENvironment (CODEEN).
Figure 1: Simplified workflow of the OU-MER pipeline. The pipeline is fed with a list of input products coming from OU-VIS, OU-NIR and OU-EXT, together with a set of configuration files and auxiliary products. The pipeline performs a mosaicing of input images before generating the source list after a combination of detection and optimal deblending procedures. The pipeline computes photometry and morphology measurements and collects all this information in a set of output catalogues. Those catalogues, together with the reference images, represent the output of the pipeline.
The pipeline is in charge of running PEs on the Euclid SGS infrastructure and controlling the parallelization scheme. The orchestration is performed via the Euclid Infrastructure Abstraction Layer (IAL), which is also in charge of managing the input/output data stream from and to the Euclid Archive System (EAS). A simplified MER PF pipeline workflow is shown in Fig. 1.
A standard run of the MER PF pipeline is performed on sky regions called MER tiles, described in Sect. 3.1, fed with the following input:
  • the definition of the MER tile to be processed (see Sect. 3.1);
  • the list of VIS calibrated frames (both nominal and short exposures, see Euclid Collaboration: McCracken et al. 2025) intersecting the tile;
  • the list of NIR calibrated frames intersecting the tile;
  • the list of EXT stacked frames covering the tile;
  • a cutout of the Gaia Data Release 3 catalogue (Gaia Collaboration et al. 2023) with the sources that fall inside the tile;
  • a set of configuration files, specifying the parameterisation of the pipeline PEs;
  • a set of ancillary files, stored in the Mission Data Base (MDB) data container;
  • the trained machine-learning models to be used for source classification;
  • a list of quality parameter ranges used to flag invalid output data products.
The MER PF outputs the following products:
  • the list of background-subtracted mosaiced images, one for each band, used to measure the source photometry;
  • the VIS and NIR background-subtracted mosaiced images used to obtain the list of detected sources;
  • the segmentation map containing the spatial pixels associated to each individual source;
  • a set of catalogues, storing all the source information measured by the pipeline.
After a first initialisation step, aimed at decompressing and sorting the input images, the MER PF pipeline produces the mosaiced images. The mosaicing step (Sect. 3.2) coadds the VIS and NIR calibrated frames. VIS and NIR stacked images, together with the EXT stacks, are then rebinned so that all the mosaiced images share the same dimensions, image centre and pixel size ($0\overset{\prime\prime}{.}1$). On each mosaic, a background subtraction algorithm (Sect. 3.3) evaluates the background of an input mosaic, subtracts it from the original mosaic, and stores the background-subtracted image in a FITS file. The goal of this step is to remove any remaining residual background, because a flat background is fundamental to obtain precise photometry measurements.
The list of source candidates is extracted from the VIS mosaic and from a stack of the three NIR bands independently (Sect. 4.1). This preliminary segmentation of the image is then refined by an optimal deblending algorithm (Sect. 4.3) in order to obtain the list of sources detected in the VIS mosaic and the one detected on the NIR stack. The two lists are then merged into one single list, taking care of source duplicates (Sect. 4.4); this represents the formal list of sources detected by the OU-MER pipeline.
At this point, the input point spread functions (PSFs) are propagated for each mosaiced image and evaluated at the detected source positions, together with the PSF convolution kernels required by the subsequent photometry steps (Sect. 5).
The list of VIS+NIR detected sources is used as a baseline to force the photometry measurements in the detected loci (Sect. 6). For each source, the MER PF pipeline measures:
  • the Kron total flux, measured on the detection mosaics;
  • PSF-matched aperture photometry on all the input bands;
  • template fitting photometry on all the input bands.
For all the sources detected in the $I_{\scriptscriptstyle\rm E}$ band, the MER PF pipeline provides a measure of the Concentration, Asymmetry, and Smoothness (CAS) parameters (Sect. 7.1), a list of morphological features measured via a machine-learning approach (Sect. 7.3) as well as a point-like probability (that can be used for star/galaxy separation, see Sect.  7.4).
A cataloguing unit then carries out the merging of all the information provided by the previous steps (Sect. 8). The main output of this step is the OU-MER final catalogue of sources with unique identification number, containing object information relevant in the Euclid project, in particular to OU-MER’s customer OUs, namely OU-PHZ, OU-SIR, and OU-SHE.
Before releasing the output products to the Euclid Archive System (EAS), a final step runs on all the mosaiced images and the final source catalogue, performing a Sérsic fit and adding further photometry and morphology information (Sect. 7.2).
Initially designed as a separate pipeline to be run on the output products of the nominal pipeline, the Sérsic fit module has been integrated rather late into the MER PF processing workflow, after extensive testing and optimization.
With the current Q1 release configuration, taking into account that many PEs run in parallel and some of them make use of more than one processing cores, the total execution time is equivalent to approximately 140 core hours. The mosaicing, background-subraction, and PSF calculation PEs account for 13% of the processing time. On average, 10% of the time is spent on the detection and deblending PEs. The large majority (75%) of the processing time is spent running the different photometry methods and on the calculation of the PSF transformation kernels.
3 Imaging input and mosaic creation
3.1 Euclid tiling
Figure 2: Example of tile geometry, specifically for EDF-N tile 10215949: centre (green), core area (magenta) and extended area (blue).
The MER PF pipeline runs on chunks of data called ‘tiles’, with each tile covering a predefined area on the sky. In every OU-MER PF run, all the available Euclid data covering the tile region are included in the processing. The Euclid tiling solution (Kümmel et al. 2022b) has a tile placement that is independent of the pattern of the Euclid observations and covers the entire sky $-88\overset{\circ}{.}0<{\rm Dec}<+88\overset{\circ}{.}0$ to become independent of the Euclid survey plan (Euclid Collaboration: Scaramella et al. 2022). The tiling solution associates each sky position with exactly one tile. The catalogues resulting from the processing of a set of adjacent tiles do not contain the same object in more than one tile. The Euclid tiling assures a certain overlap between two adjacent tiles to guarantee that the tile boundaries do not interfere with computation of the objects’ properties.
In the Euclid tiling solution each tile is defined by its central position in $({\rm RA,Dec})$ and size (width, height), and identified by a unique tile identifier. Every tile has an extended area and a core area. While the extended areas of adjacent tiles have an overlap to avoid border problems in the measurements, the core areas of any two tiles do not overlap. The extended areas and the core areas of all tiles cover the entire sky, with the limits mentioned above. The imaging output of OU-MER covers the extended tile area and object detection and property measurements are done on the extended tile area, however, the catalogue output is delivered only for the core area, thus avoiding multiple outputs of the same object. The core areas of the tiles are defined as HEALPix indices (Górski et al. 2005). Figure 2 shows tile 10215949 in Euclid Deep Field North (EDF-N), with $\rm(RA,Dec)=(272\overset{\circ}{.}9038602,66\overset{\circ}{.}5)$ and its centre (green), extended area (blue), and the HEALPix pixels (magenta) defining the core area.
Figure 3: Example tiling pattern. Notice that, the height of tile 102158273 in the EDF-N is enlarged to fully cover a larger galaxy.
The Euclid tiling solution:
  • has an extended tile area of $32^{\prime}\times 32^{\prime}$ in tangential projection;
  • increments the tile centers by $30^{\prime}$, which leads to an overlap of $2^{\prime}$ for adjacent tiles;
  • starts the central tile positions at $\rm(RA,Dec)=(0\overset{\circ}{.}0,0\overset{\circ}{.}0)$ and places them in strips of constant Dec;
  • adjusts the tile width when closing a strip at the Meridian;
  • defines the core area using HEALPix indices with $N_{\text{side}}=13$;
To define the core areas we first start, for each tile, with the HEALPix indices that are inside of the extended area. Then we identify the indices located in the overlap of extended areas, which appear in more than one tile. For those, we measure the mininum distance to the border of each tile in which they appear; then each index is attached to the tile which maximizes this distance and removed from all others. This procedure avoids duplicated HEALPix indices in several tile an assures a contiguous coverage of the core areas.
With a strict tiling as outlined so far, some of the galaxies with sizes larger than the overlap size will inevitably extend beyond the tile that its centre is located on. To solve these cases, we identified, from publicly available catalogues, all large galaxies ($\rm size>2^{\prime}$) located in the Q1 area and store their position and size. We identify the critical cases where a large galaxy extends between more than one tile. We allow the tile with the galaxy centre to ‘breathe’, which means to enlarge its size until it contains the entire galaxy. This procedure covered all large galaxies in Q1. Figure 3 shows the tiling in the lower part of the EDF-N. In this particular case, the height of tile 102158273 was enlarged to $33^{\prime}$ in order to fully cover the galaxy PGC61126.
3.2 Mosaicing
For the later measurement steps, specifically the photometry with A-PHOT and T-PHOT (see Sect. 6), we need a multi-band set of co-aligned images, where each pixel represent the same position on the sky. The layout of these co-aligned images is given by setting a pixel scale of $0\overset{\prime\prime}{.}1\,{\rm pixel}^{-1}$ (to properly sample the VIS data) and the dimensions of the Euclid tiles presented in Sect. 3.1.
EXT imaging data is already provided per tile as stacked frames in the native scale of the respective camera, e.g. $0\overset{\prime\prime}{.}264\,{\rm pixel}^{-1}$ for the DES data. These data are rebinned to the common pixel scale using a BILINEAR interpolation kernel.
VIS and NIR input data are the set of fully calibrated frames covering a tile in their native pixel scale of $0\overset{\prime\prime}{.}1\,{\rm pixel}^{-1}$ and $0\overset{\prime\prime}{.}3\,{\rm pixel}^{-1}$, respectively. After subtracting the background estimate provided with the input data (see Euclid Collaboration: McCracken et al. 2025; Euclid Collaboration: Polenta et al. 2025), these calibrated frames are rebinned and stacked to deep co-added images with a BILINEAR interpolation kernel and the combine type MEDIAN. For the stacking we apply the RMS data provided with the VIS and NIR data as weights. The different zero points for the VIS and NIR data are translated to global weighting values which are used in the co-add process. In addition, we scale these global weights such that the zero point of the resulting co-added images corresponds to the median value of the input zero points, rounded to one decimal place.
For all rebinning and co-adding operations we use the software CT_SWarp, which is a version of the SWarp (Bertin et al. 2002) software that is fully integrated into the Euclid processing environment.
In CT_SWarp we also implemented the interpolation kernel BILINEAR_FLAGS, which allows to combine the flagging information for the VIS and NIR data with the bitwise XOR operation. The resulting flag image shows, for every pixel, the flag values that were associated to the corresponding input pixels, which usually resulted in the exclusion of the flagged pixels.
The result of this processing step is, for every filter band, the set of co-added image plus the corresponding RMS and flag images. The astrometric validation of the co-added images shows that the non-linear astrometric solutions provided with the VIS and NIR data fulfill the requirements (see Sect. 9.1).
3.3 Background subtraction
We run the background subtraction step to remove possible residual background from the mosaic images. To avoid contamination from bright extended sources in the input mosaic, we apply a segmentation mask generated by source detection run using SourceExtractor++ (SE++, Bertin et al. 2020; Kümmel et al. 2022a) which is a re-designed and extended version of the SExtractor2 (SE2, Bertin & Arnouts 1996) software. The masked mosaic is divided into a grid, for each cell of which we estimate the background as a combination of k-sigma clipping and median filtering. In each cell, the local background pixel distribution is clipped iteratively until convergence at $2\sigma$ around its mean. If after convergence the mean and the median background differ by more than $3\sigma$ (e.g., in crowded fields), the background on that cell is defined to be the mean of the clipped histogram. If the mean and the median differ by less than $3\sigma$, the background on that cell is estimated as the mode of the distribution,
$B=2.5\,M-1.5\,\mu,$ (1)
where $M$ and $\mu$ are the median and mean of the clipped histogram, respectively.
Eventual gaps (cells lacking unflagged pixels to estimate the background) are filled with the distance-weighted average of the background values of the cells at the edges of the gap. Once the background is computed in each cell of the grid, we apply a median filter to suppress possible local overestimations due to residuals from bright stars. The background model at the original resolution of the mosaic is determined by a spline-interpolation of the cell values, and is subtracted from the mosaic image. Two main parameters control the background generation: the grid size, which we set to $256\times 256$ pixels, and the size of the filter box, which we set to $3$ cells.
4 Assembly of the source list
4.1 General considerations for source detection and deblending
The strategy for detecting objects in the available imaging data is dictated by the Euclid project requirements:
  • Objects shall only be detected in the photometric bands coming directly from the Euclid instruments, which are $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$. VIS and NIR imaging data for a specific tile are observed over a short time (around $1\,\rm{h}$) and give consistent basis for object detection. EXT data can span a large range of epochs, even within the bands available for one tile. Due to the epoch differences this would result in astrometric offsets within the data and pose problems for the photometry.
  • We need to detect all objects in the Euclid VIS band ($I_{\scriptscriptstyle\rm E}$). The weak lensing measurements by OU-SHE require an object sample with a simple selection function. Deriving the weak lensing sample from objects detected outside of VIS, such as co-added or $\chi^{2}$ images in $I_{\scriptscriptstyle\rm E}$ +$Y_{\scriptscriptstyle\rm E}$/$J_{\scriptscriptstyle\rm E}$/$H_{\scriptscriptstyle\rm E}$ would result in biases and systematic effects that cannot be recovered in the shear analysis.
  • We need to detect all objects in the Euclid NIR data ($Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, $H_{\scriptscriptstyle\rm E}$). The catalogues from the MER PF are used as input to the slitless spectroscopy analysis. This requires a deep and complete detection process to find the direct image counterparts of all objects that possibly have spectral traces or emission lines in the slitless spectroscopic NISP data and hence must include a detection in the NIR bands.
  • The detection shall record whether a source was detected in VIS or in NIR.
Since we need to separate sources detected in VIS and in NIR we cannot work with a single, deep detection image such as VIS+NIR or a $\chi^{2}$ image as described in Szalay et al. (1999). Instead, as outlined in Fig.  1, we take the following the approach:
  • perform a source detection and deblending in VIS.
  • perform an independent source detection and deblending in NIR.
  • Combine the VIS sources and the NIR sources to a common object list and segmentation map.
  • The computation of object properties, such as photometry and morphology, is then performed on the combined object lists.
4.2 Source detection
Figure 4: Small cutouts of the detection images $I_{\scriptscriptstyle\rm E}$ (left) and NIR ($Y_{\scriptscriptstyle\rm E}+J_{\scriptscriptstyle\rm E}+H_{\scriptscriptstyle\rm E}$, right) in tile 102021017. The VIS and NIR detected sources are marked in the left and right panels, respectively
The source detection is done using SE++. The typical procedure for detecting objects in astronomical images is to:
  • convolve the data and the rms images with the segmentation filter in order to enhance the signal-to-noise ratio (S/N);
[…]
With this method there are usually some false detections around large, extended galaxies. Using reasonable values for the parameters detection-threshold and detection-minimum-area, which are tailored to avoid false detections from noise peaks in the background, the detection cannot follow the extended objects far out into the sky background. The algorithm then picks up ‘objects’ in the extended parts of large galaxies as individual sources.
To avoid these false positive detections, we have implemented an improved detection scheme in SE++ that reduces the number of false detections around bright galaxies. In this detection scheme there is in addition to the detection-threshold and detection-minimum-area discussed above a second threshold with:
  • core-threshold-value: the SNR threshold value;
  • core-minimum-area: the number of pixels above the core-threshold-value an objects needs to have to validate the object;
With these two detection thresholds it is possible to have a rather low value for the detection-threshold parameter. Then the detection areas of large galaxies extend into the region of the fake objects. The second, higher threshold core-threshold-value then eliminates the false detections that result from the low value detection-threshold at noise peaks in the sky background. The lower value of the detection-threshold in the ‘deep detection’ leads to a larger detection area (around) of the objects.
The VIS detection image is the background-subtracted mosaiced image described in Sects. 3.2 and 3.3. For the detection in NIR we produce a deep NIR detection image from the $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ background-subtracted mosaiced images using weighted summation. The different zero points of the input images are taken into account in the weighted summation to yield a common zero point for the resulting deep NIR detection image.
Table 1: Important SE++ parameters used for the VIS and NIR detections.
Parameter VIS NIR
segmentation-filter (Gaussian FWHM [pix]) $2.4$ $5.0$
detection-threshold 0.9 0.4
detection-minimum-area 10 10
core-threshold-value 1.3 0.55
core-minimum-area 10 10
Table 1 lists the most important parameters used in the VIS and NIR detection. The segmentation filters are chosen to have a width similar to the target population, which are faint galaxies. The other parameters have been ‘calibrated’ using simulated images to produce a rate of false detections $<1\%$, which is a typical value for large-scale surveys.
4.3 Deblending
Deblending is an essential procedure for separating overlapping sources in astronomical images. With its unprecedented sensitivity and resolution, Euclid captures densely packed fields and intricate morphologies, making source disentanglement one of the key challenges in its data analysis.
Following the independent image segmentation in the VIS mosaic and the NIR stack, resulting from detection, deblending is executed in a parallel manner for both groups of source clusters. The aim is to put together a refined list of individual sources in the merging stage described in the next section.
The deblending process is carried out using a dedicated software component based on ASTErIsM (Tramacere et al. 2016). ASTErIsM is a Python package designed for automatic source detection and classification, leveraging two topometric clustering algorithms: DBSCAN (Ester et al. 1996) for identifying sources and DENCLUE (Hinneburg et al. 1998; Hinneburg & Gabriel 2007) for deblending overlapping sources and identifying galaxy morphological features such as spiral arms. Our internal processing element, with a specific set of configuration parameters for VIS mosaic and NIR stack, inherits the operation of ASTErIsM as a deblender, which is based on the execution of a modified version of the DENCLUE density-based clustering algorithm, where the kernel-density estimation is substituted with a convolution of the image using a given kernel. This algorithm processes, as a function of the configuration parameters, a downsampled version of the segmentation clusters in terms of a distinct spatial scale. The adaptive approach allows for a robust separation of the overlapping sources within each cluster.
The two sets of parameters selected in the MER pipeline for deblending the clusters of sources detected in the VIS mosaic and the NIR stack have been validated using the images and catalogues from the True Universe (TU) simulation (Euclid Collaboration: Serrano et al. 2024). This was done by exploring a discrete sampling of the parameter space to single out the values for which the total number of over- and under-deblendings is minimised in reference to a ground-truth segmentation map built from the TU catalogue. Two parameters were considered in this process due to their significant influence on the operation of the DENCLUE algorithm, one of them defines the typical scale of the sources blended within a cluster, while the other one establishes the converge criterion that identifies the local maximum corresponding to the centre of a deblended source.
4.4 Merging the VIS and NIR detections
The results of the independent detection and deblending in VIS and NIR are one catalogue and one segmentation image in each of the VIS and NIR bands. These two data sets need to be combined to a single output catalogue and the corresponding segmentation image for the subsequent determination of object properties (photometry, morphology, etc.).
The VIS data have higher resolution ($0\overset{\prime\prime}{.}1$) than NIR ($0\overset{\prime\prime}{.}3$) and a better sampling. This leads to much better results for the detection and deblending. As a consequence, we combine the VIS and NIR datasets such that the merged catalogue and segmentation image contain all VIS objects plus the so-called NIR-only objects, which are the objects that are detect in NIR but not in VIS.
To identify the NIR-only objects we project the coordinates of all NIR detected objects into the VIS segmentation image. NIR objects that are already detected in VIS are identified by evaluating the VIS segmentation image, which marks the pixels of all VIS detectd sources. VIS segmentation values ¿ 0 at the projected central NIR position are considered to be already detected in VIS, since they are already part of a VIS object. NIR objects with value 0 in the VIS segmentation image are not detected in VIS and thus form the subset of NIR-only sources. The final output catalogue is then the combined VIS and NIR-only catalogue.
To derive the combined VIS+NIR segmentation image we copy the VIS segmentation image and add the detected pixels of the NIR-only sources from the NIR segmentation image to the now combined segmentation image. In this operation we do not overwrite any VIS-detected pixels. Preserving all VIS-detected pixels changes and reduces the detection area of few NIR-only objects, and we recompute their properties (position and shape). Also we discard NIR-only objects that would end up having fewer than detection-minimum-area pixels (see Tab. 1) after the combination of the segmentation maps.
Figure 4 shows in the left and right panel a small cutout from the VIS and NIR detection images, respectively. The red circles mark all sources detected on these cutouts. The brighter objects are all detected in VIS and the NIR detected sources constitute the faint and rather red sources. There are $141$ and $31$ sources detected on the VIS and NIR cutout, respectively. As in this small cutout about $20\%$ of all sources are detected in NIR.
It can sometimes happen in the combined VIS+NIR segmentation map that VIS and NIR-only sources are located side-by-side without any empty sky pixels in between. These artificial VIS-NIR blends are identified and marked by assigning a unique number to the corresponding objects in the final catalogue column PARENT_VISNIR.
We are aware that current merging procedure may dis-favor very interesting sources such as red objects that are very close to VIS objects. Some of those sources should be discovered in dedicated pipelines such as The Strong Lensing Discovery Engine (see Euclid Collaboration: Walmsley et al. 2025a), but we are also planning to refine the merging procedure in the future.
5 Mapping of the point spread functions and convolution kernels
MER receives the PSFs measured by the VIS, NIR, and EXT PFs and propagates them to the MER mosaic grid to calculate a different PSF stamp for each individual source and band. VIS PSFs are measured for each VIS quadrant on a dedicated calibration processing run using the PSFEx software (Bertin 2011) and are assumed to remain constant for each VIS calibrated frame observation (Euclid Collaboration: McCracken et al. 2025). NIR PSFs are calculated for each NISP detector and each calibrated frame (Euclid Collaboration: Polenta et al. 2025) using PSFEx.
MER propagates the VIS and NIR PSFs following a similar approach as described in Sect. 3.2, with the following differences:
  • the pixel values of the mosaicing step input images are set to zero, but the zero points, WCS information, flags, and RMS are left untouched.
  • PSF stamps are then added to the empty images at the deblended source positions, taking into account the VIS and NIR PSF spatial variations and the image zero points.
  • the Swarp resampling and co-adding software (Bertin et al. 2002) is run on the modified images and a mosaic with the averaged stamp PSFs is produced.
  • the averaged PSF stamps are extracted for each source and stored in a FITS file, which contains a grid with the PSF stamps and a table with their associated sky and pixel coordinates.
Source separations in the MER catalogue can be smaller than the input VIS and NIR PSF stamp sizes (less than 33″3  ″). To avoid any overlap between the PSF stamps of nearby sources the whole process is repeated three times. Each run uses a different subset of sources whose separations are large enough to avoid PSF stamp overlaps. This generally produces a PSF stamp for more than 80% of the sources. For the remaining sources, the closest PSF stamp is used, which by design is at a distance smaller than the PSF stamp size.
In its current state, the MER pipeline assumes that the PSF of a given band does not change with the source intrinsic color. The propagation of wavelength-dependent PSFs could be introduced in the future if the VIS and NIR PFs are able to deliver their input PSFs with this information.
The EXT stage-2 PF calculates their stacked frames PSFs using a similar approach as described above. MER rebins the EXT PSFs from the EXT stacked frame pixel size to the reference $0.1″$ mosaic pixel size. The PSF stamp positions are based on the EXT stacked frame catalogue, which is generally not as deep as the MER catalogue, and as a result contains fewer sources. We select the closest PSF stamp for each source in the MER catalogue.
Table 2: Full width half maximum (FWHM) values measured using two-dimensional Gaussian fits on the MER propagated PSFs and the Gaia star stamps extracted from the MER mosaics. The different statistics are calculated combining the median FWHM values measured for each tile in the Q1 data release. Tiles with less than 20% coverage in one of the Euclid bands, as well as tiles associated to the Dark Cloud region, have not been considered.
Survey Band PSF FWHM Gaia stars FWHM
Median Min Max Median Min Max
EWS $I_{\scriptscriptstyle\rm E}$ 0.203 0.197 0.212 0.204 0.199 0.213
EWS $Y_{\scriptscriptstyle\rm E}$ 0.475 0.456 0.484 0.492 0.474 0.502
EWS $J_{\scriptscriptstyle\rm E}$ 0.504 0.494 0.517 0.513 0.505 0.524
EWS $H_{\scriptscriptstyle\rm E}$ 0.542 0.536 0.547 0.552 0.539 0.570
UNIONS CFIS $u$ 1.00 0.84 1.16 0.99 0.84 1.17
UNIONS WHIGS $g$ 0.79 0.57 1.47 0.79 0.58 1.48
UNIONS CFIS $r$ 0.85 0.69 1.00 0.85 0.70 0.99
UNIONS Pan-STARRS $i$ 1.27 1.18 1.37 1.26 1.17 1.36
UNIONS WISHES $z$ 0.63 0.51 0.83 0.63 0.51 0.83
DES $g$ 1.28 1.02 1.40 1.28 1.01 1.43
DES $r$ 1.12 0.95 1.23 1.12 0.95 1.26
DES $i$ 1.07 1.00 1.47 1.07 0.99 1.47
DES $z$ 1.05 0.95 1.27 1.06 0.95 1.27
Table 2 compares the typical Full Width at Half Maximum (FWHM) values measured on the propagated PSF stamps and the detected Gaia stars in the MER mosaics. The agreement between the median values is better than 1%, with the exception of the NISP bands. The propagated PSFs for these bands are between 1.75% and 3.5% more compact than the Gaia star profiles. Future versions of the NIR PF pipeline should improve the quality of the PSFs that MER receives as input. We expect that this will reduce the FWHM discrepancies that we currently find in the Q1 processing.
The FWHM values reported for the propagated VIS and NIR PSFs in Table 2 are 30% to 35% larger than the corresponding input PSFs measured on the VIS and NIR calibrated frames (Euclid Collaboration: McCracken et al. 2025; Euclid Collaboration: Polenta et al. 2025). This is a result of the stacking procedure described in Sect. 3.2. In particular, the use of a BILINEAR interpolation kernel instead of a LANCZOS3 kernel introduces an extra 15% broadening of the PSFs.
The multi-band photometry (described in Sect. 6) needs convolution kernels to homogenise the PSFs from the different bands. The algorithm we use to create the convolution kernels is based on Wiener filtering with a tunable regularization parameter. For each detected object, we compute convolution kernels from higher- to lower-resolution bands using the object PSF stamps generated in the previous step. A complete description of the kernel creation algorithm can be found in Boucaud et al. (2016).
6 Multi-band photometry
OU-MER must provide photometric measurements for all the detected sources. The pipeline has been designed to produce photometric estimates using various techniques, which can be combined in different ways to obtain the desired scientific information. Namely, the output catalogue contains: (i) aperture photometry, obtained using a customized version of the software A-PHOT (Merlin et al. 2019); (ii) template-fitting photometry, obtained using a customized version of the code T-PHOT (Merlin et al. 2015, 2016); and (iii) model-fitting photometry, obtained using SE++.
Concerning aperture photometry, the total flux of each object is estimated as the flux within a Kron (1980) elliptical aperture on the detection image, which is the $I_{\scriptscriptstyle\rm E}$ mosaic for most of the sources, and the NIR stack for the additional NIR-detected sources (see Sect. 3.2). This corresponds to the FLUX_AUTO in SE2; here the Kron radius is computed as
$R_{\rm Kron}=\max\left(7\,a,\frac{\sum_{ij}r_{ij}\,f(i,j)}{\sum_{ij}f(i,j)}\right),$ (2)
where $a$ is the semi-major axis provided by the deblending algorithm during the deblending step, $r_{ij}$ is the distance of each pixel to the centroid of the object, $f(i,j)$ is its flux, and the summation is extended over a circular region typically of radius $6\,a$. The final Kron area for which the flux is computed is an elliptical area of semi-major axis $2.5\,R_{\rm Kron}$ and ellipticity, again provided by the deblending step. A-PHOT sub-divides the pixels crossed by the limit of the elliptical aperture into 10 subpixels, to provide a more precise estimate.
Aperture photometry is then measured for all bands on PSF-matched images, within a set of circular apertures with diameters fixed individually source by source as integer multiples of the FWHM of the worst resolution band (typically an EXT band, with FWHM around $1^{\prime\prime}$). In this way, the same physical region of the sources is always sampled, providing a robust colour estimate. The smoothing is performed using convolution kernels obtained from the PSFs estimated as described in Sect. 5. To determine the total flux in each band, one must combine the value obtained on the detection band with these aperture fluxes, as follows (e.g. Merlin et al. 2022):
$f_{\rm tot,band}=f_{\rm ap,band}\,\frac{f_{\rm tot,det}}{f_{\rm ap,det}}.$ (3)
The uncertainty budget can be computed with an analogous formula, where it is safe to assume that the ratio $f_{\rm tot,det}/f_{\rm ap,det}$ is a constant and there is no need to propagate their errors. The uncertainty $e_{\rm ap,band}$ is computed by A-PHOT as the square root of the quadratic sum of the pixels in the PSF-matched RMS map, within the considered aperture. All the measurements are done without performing a further local background subtraction, since two background subtraction steps have already been performed on the images.
Template-fitting photometry is a well-established technique used to exploit spatial and morphological information from high-resolution images to estimate fluxes on lower-resolution images of the same FoV, using cut-outs of sources from the former (smoothed with a convolution kernel) as priors for the latter, to mitigate the effect of blending of sources. While it was initially designed to fit infrared data (from e.g., Spitzer) with HST priors, it has proven to be useful in the Euclid context as well. The crispness and depth of the $I_{\scriptscriptstyle\rm E}$ data can provide exquisite priors for EXT data, and can be used for NIR data as well, although in that case the difference in resolution is not dramatic. Fluxes are obtained by solving a $\chi^{2}$ minimisation problem,
$\chi^{2}=\sum_{m,n}\left[\frac{I(m,n)-\sum_{i}^{N}F_{i}(m,n)P_{i}(m,n)}{\sigma(m,n)}\right]^{2},$ (4)
where $m$ and $n$ are pixel indexes, $N$ is the number of sources being fitted, $P$ are the values of the pixels of the normalised high-resolution cutouts convolved to the measurement image resolution, $I$ are the values of the pixels in the measurement image, and $\sigma$ are the values of the pixels of its RMS map (see Merlin et al. 2015, for details). The error budget of the fitting process is obtained from the covariance matrix of the problem, as the square root of the diagonal element relative to that source (i.e. its variance). In principle, one could fit the full image at once, but in practice this would be too computationally demanding given the large spatial dimensions of the OU-MER tiles. The pipeline therefore exploits a feature of T-PHOT which allows to perform individual fits looping on the sources, building a cell around each one including its contaminants, and removing if after it has been fitted. This algorithm has proven to yield a significant saving of computational time, while keeping optimal accuracy.
The pipeline provides the direct output of the template-fitting runs, which can be taken as a proxy for the total flux of the objects. Additionally, fluxes measured in the detection band smoothed to the resolution of each other band are also provided; they can be used to obtain an estimate of colours complementary to that obtained with aperture photometry, and used to scale the detection Kron flux to get another total flux value in each band. Typically, one expects $f_{\rm templ,band}\leq f_{\rm tot,band}$, where the latter is the total flux from aperture photometry as described above; this is because the contamination from nearby sources, when present, should be drastically decreased. Figure 5 shows the situation for a random OU-MER tile, and the case of one discrepant galaxy. However, the colors obtained with the two techniques should obviously be reasonably consistent. Figure 6 shows the compared $g$ $H_{\scriptscriptstyle\rm E}$ colors ($g$ is from DECAM) obtained from aperture and template fitting photometry in the same tile, for stars (blue crosses, POINT_LIKE_PROB$>0.7$) and galaxies (red dots, POINT_LIKE_PROB0.7absent0.7$\leq$ 0.7≤ 0.7; see Sect. 7.4), ensuring this is indeed the case. It should be stressed that for Q1 we found that for point-like sources (bright stars in particular), contrary to what was just discussed for extended sources, the total fluxes computed using aperture photometry were found to be more accurate than template-fitting fluxes in NIR bands, most likely due to their higher sensitivity to PSF related issues (see Sect. 9.5). Currently, OU-PHZ uses aperture colors to compute photometric redshifts (see Euclid Collaboration: Tucci et al. 2025).
Figure 5: Upper panel: difference in $H_{\scriptscriptstyle\rm E}$ total magnitudes as obtained with template-fitting and aperture photometry as a function of $I_{\scriptscriptstyle\rm E}$ detection magnitude, for extended objects (POINT_LIKE_PROB$<0.7$, see Sect. 7.4) in a random OU-MER tile. Template-fitting magnitudes are typically fainter, because the contaminating light from blended sources is removed in the fitting process. Lower panel: as an example, the source selected with the black circle in the upper panel is shown in the corresponding $H_{\scriptscriptstyle\rm E}$ OU-MER mosaic. The AB magnitude of the pixels within the white aperture (which is certainly still contaminated by the nearby objects), as given by the SAO ds9 software, is 19.96; the OU-MER catalog values are 19.81$\pm$0.02 (total from color in 2FWHM aperture) and 19.99±$\pm 0.01$ (template-fitting).
Figure 6: Difference in $g$-$H_{\scriptscriptstyle\rm E}$ colors obtained from aperture and template fitting photomety for a random OU-MER tile, as a function of $I_{\scriptscriptstyle\rm E}$ magnitude.
We point out that template-fitting, where real cut-outs of high-resolution images are used to infer the expected shape of low-resolution images, is different from model-fitting, where galaxy profiles are fitted using analytical models. Indeed, the pipeline also provides photometric values derived by fitting a parameterized model to the multi-band imaging data. For Q1 we have chosen to fit Sérsic models (Sérsic 1963) to all detected objects. This is done with SE++ in an iterative procedure that involves VIS and NIR data to fix the Sérsic parameters and then EXT data for photometry only. Further details are given in Sect. 7.2.
It is worth recalling that, as it is custom practice, the provided error budgets are a measure of the uncertainty due to the observational noise level, as encoded in the RMS map. In other words, it does not include an estimate of any systematic deviation to which a measurement might be prone (e.g. because of contamination from nearby sources), and might therefore be an underestimation of the actual uncertainty on the real flux of the sources.
The flux within the segmented area of each source is also provided (corresponding to FLUX_ISO in SE2). Finally, a PSF-fitting estimate is provided for the VIS band only, again using T-PHOT.
All the photometric measurements are stored in the MER photometric catalogues (Sect. 8), a column-by-column description of which is referenced in Appendix A. It is not always trivial, though, to handle the amount of information stored in such a huge container. In order to help the user to become familiar with our products and to handle the photometric information measured by OU-MER, we provide a Photometry Cookbook [^1] http://st-dm.pages.euclid-sgs.uk/-/data-product-doc/-/jobs/162527/artifacts/build/merdpd/merphotometrycookbook.html , explaining the basic operations that need to be performed in order to correctly use MER data.
7 Morphological characterization
For each galaxy used in the weak lensing analysis up to the limiting magnitude the following information is provided: (i) the classification star/galaxy (ii) at least one size measurement. An object radius is provided via the detection and photometry pipelines. Three kinds of morphological estimators are included in the MER catalogue:
  • Non-parametric parameters. Those are Concentration, Asymmetry, and Smoothness (CAS, Conselice 2014; Tohill et al. 2021), the Gini index (Lotz et al. 2004), and the second order moment at 20% flux, i.e the M20 index;
  • Parametric parameters computed through model fitting (Euclid Collaboration: Bretonnière et al. 2022, 2023);
  • Deep learning based morphologies calibrated on visual classifications.
As per the photometric catalogue, we provide a Morphology Cookbook [^2] http://st-dm.pages.euclid-sgs.uk/-/data-product-doc/-/jobs/162527/artifacts/build/merdpd/mermorphologycookbook.html aimed at better understanding how to correctly use the information stored in the MER morphology catalogue.
7.1 Non-parametric indices for object morphology
The original implementation of CAS, Gini, and M20 parameters as Euclid processing element was coded from ASTErIsM, a code for deblending objects and morphological measurements (Tramacere et al. 2016). The CAS parameters are non-parametric parameters carrying morphological characteristic in a three-dimensional space that segregate main morphological features. They are computed on Euclid $I_{\scriptscriptstyle\rm E}$ images covering the segmentation map of each object.
7.1.1 Concentration index $C$
The concentration of light index $C$ (Conselice 2014) is used as a method for quantifying how much light is in the centre of a galaxy as opposed to its outer parts. It correlates strongly with Sérsic $n$ values,
$C:=5\log_{10}\frac{r_{\text{80}}}{r_{\text{20}}},$ (5)
where $r_{\text{80}},r_{\text{20}}$ are the circular apertures enclosing $80\%$ and $20\%$ of the total galaxy flux. The associated error is defined:
$C_{\mathrm{err}}:=5\operatorname{log_{10}}\frac{r_{80}}{r_{20}}-5\operatorname{log_{10}}\frac{r_{80}-0.5}{r_{20}+0.5}.$ (6)
Figure 7: Concentration calculation schematics, here $r_{80}$ and $r_{20}$ are circular radii encompassing 80% and 20% of the galaxy flux (cf text and Eq. 5).
7.1.2 Asymmetry $A$
The asymmetry index $A$ (Conselice 2014) measures how asymmetric a galaxy is after rotating along the line of sight centre axis of the galaxy by $180.0°$ (Fig. 8). The analytic definition calculating $A$ is given by
$A:=\min\left(\frac{\sum|I_{0}-I_{180}|}{\sum|I_{0}|}\right)-\min\left(\frac{\sum|B_{0}-B_{180}|}{\sum|I_{0}|}\right),$ (7)
[…]
$A_{\mathrm{err}}:=\frac{\sigma[B(x,y)]}{f+\sqrt{|f|}}.$ (8)
[…]
7.1.3 Clumpiness/Smoothness index $S$
The clumpiness (or smoothness) index S (Conselice 2014) is used to describe the fraction of light in a galaxy that is contained in clumps. The smoothness can by computed as
$S:=\frac{f_{s}-f_{b}}{f},$ (9)
with
$f_{s}$ ::=:= $\sum I(x,y)-I(x,y)_{\mathrm{blurr}},$ (10)
$f_{b}$ ::=:= $\sum B(x,y)-B(x,y)_{\mathrm{blurr}},$ (11)
$f$ ::=:= $\sum I(x,y),$ (12)
where $I(x,y)$ are the image pixels chopped within $r_{\rm inner}=0.1\eta$ of the object flux peak, $\eta$ is the petrosian radius (Petrosian, 1976), $I(x,y)_{\mathrm{blurr}}$ is the chopped image convolved with a gaussian kernel of $\sigma=0.1\eta$, $B(x,y)$, and $B(x,y)_{\mathrm{blurr}}$ are the noise maps treated identically. Finally, the associated error is
$S_{\mathrm{err}}:=\frac{10\sqrt{|nf_{s}|}}{n(f+\sqrt{|f|})},$ (13)
where $n$ is the number of unmasked pixels.
7.1.4 Gini index
The Gini index (Lotz et al. 2004) originates from the world of economics, where it has been defined originally to evaluate inequalities of incomes in a population. If everyone earns the same income, Gini = 0, if one person concentrate all incomes Gini = 1. Applied on an astronomical image, the calculation is done on pixels covered by the segmentation map of objects. A galaxy in this case is considered a system with $n$ pixels each with a flux $f_{i}$, where $i$ ranges from 1 to $n$. The Gini index is then measured by
$\text{Gini}:=\frac{1}{|\bar{f}|n(n-1)}\sum_{i}^{n}(2i-n-1)|f_{i}|\,,$ (14)
where $\bar{f}$ is the average pixel flux value.
7.1.5 Moment-20 index M20
The mathematical definition of the M20 index is
[…]
The value of $M_{\text{tot}}$ is
$M_{\text{tot}}=\sum_{i}^{n}M_{i}=\sum_{i}^{n}f_{i}[(x_{i}-x_{c})^{2}+(y_{i}-y_{c})^{2}],$ (16)
where ($x_{c}$,$y_{c}$) is the flux-weighted mean position of the galaxy. In the case of M20, this center is defined as the location where the value of $M_{\text{tot}}$ is minimal.
7.2 Parametric morphology
To derive some parametric morphology, we have decided to fit two-dimensional Sérsic profiles (Sérsic 1963) in all photometric bands to all objects. As a function of the angular radial $r$, the Sérsic function defines the variation of the light intensity as:
$I(r)=I_{\mathrm{e}}\exp\left\{-b_{n}\left[\left(\frac{r}{R_{\mathrm{e}}}\right)^{1/n}-1\right]\right\},$ (17)
with $R_{\mathrm{e}}$ the major-axis of the elliptical profile that encloses half of the total light, $I_{\mathrm{e}}=I(R_{\mathrm{e}})$ the light intensity at $R_{\mathrm{e}}$, $n$ the Sérsic index characterizing the steepness of the profile, and $b_{n}$ a normalisation parameter which depends solely on $n$. Fitting 2D elliptical light profiles to the Euclid imaging data adds the position angle and the axis ratio.
The Sérsic fitting is performed using SE++ (Kümmel et al. 2022a), which, from version 0.21333https://github.com/astrorama/SourceXtractorPlusPlus on, includes model fitting as an isolated process without the preceding object detection process. All necessary parameters such as objects centres, the fitting area, and initial values for the model parameters are provided in an input table to the model fitting process.
The model fitting in the OU-MER PF has a dual purpose and provides, in addition to the Sérsic parameters, also the Sérsic flux in all photometric bands as an independent photometry in addition the ones introduced in Sect. 6. We do this in an iterative process:
  • The first model fitting operation on VIS and NIR data computes all structural Sérsic parameters and the VIS and NIR fluxes:
    • We fit to the VIS ($I_{\scriptscriptstyle\rm E}$) data the corresponding Sérsic radius ($R_{\rm VIS}$), ellipticity ($e_{\rm VIS}$), and index ($n_{\rm VIS}$);
    • We also fit to the NIR ($Y_{\scriptscriptstyle\rm E}$/$J_{\scriptscriptstyle\rm E}$/$H_{\scriptscriptstyle\rm E}$) data the independent Sérsic radius ($R_{\rm NIR}$), ellipticity ($e_{\rm NIR}$), and index ($n_{\rm NIR}$) for the NIR bands.
    • The VIS and NIR Sérsic models share the same angle $\alpha_{\rm VIS,NIR}$
[…]
  • The second model fitting operation runs on the EXT data $g,r,i,z$. It uses the Sérsic parameters $R_{\rm VIS}$, $e_{\rm VIS}$, $n_{\rm VIS}$ and $\alpha_{\rm VIS,NIR}$ as constant values to compute the integrated Sérsic flux in those EXT bands.
  • In both fitting operations we use the object centres provided as an input as the fixed centres for the Sérsic models.
For all fitted quantities we provide the error obtained from the square root of the diagonal element of the covariance matrix. The approach of using an iterative approach for VIS plus NIR for the morphology and then EXT for photometry is rather unusual, and configuration for deriving two independent set of Sérsic parameters with a common orientation angle is complex. As an example for this setup we share a multi-band data set [^4] https://cloud.physik.lmu.de/index.php/s/3K4KemBsw5y9yqd. As imaging data we provide cutouts from a Q1 tile. There are also the various configuration files for both model fitting operations described above and the SE++ commands to run the SE++ model fitting on the data. A scientific evaluation of the parametric morphology provided in the OU-MER catalogues is given in Euclid Collaboration: Quilley et al. (2025).
7.3 Machine learning based morphology - Zoobot
The MER catalogue contains visual morphologies following the GalaxyZoo (GZ) classification tree. This includes broad morphological classes, such as ‘featured’ or ‘smooth’, as well as more detailed features such as the count and tightness of spiral arms or the strength of stellar bars. For the Q1 data release, the MER catalogue includes visual morphologies for galaxies brighter than $I_{\scriptscriptstyle\rm E}$ $<20.5$ or with a segmentation area larger than 1 20012001\,2001 200 pixels, which represent around $1\%$ of galaxies in the photometric catalogue. This conservative selection is performed to guarantee that the structure is unambiguously detected but could be extended to fainter and smaller sources in future releases.
[…]
To optimise the classifications for Euclid, we fine-tune Zoobot with an additional 1.56 million labels obtained on Euclid stamps through a dedicated GZ campaign carried out during August 2024. Specific details about the methodology and reliability of the classifications are presented in Euclid Collaboration: Walmsley et al. (2025b).
7.4 Star-galaxy separation
A simple star-galaxy (S/G) classifier has been historically implemented in the MER pipeline in order to identify the point-like detected objects on which the PSF characterization could be performed by SHE. In addition to the POINT_LIKE_FLAG, a point-like probability (POINT_LIKE_PROB) was also requested in the output catalogue. Note that this classifier is heavily biased towards a high purity, and thus has a low completeness.
The method is inspired by the SPREAD_MODEL method provided by SE2 which is used in Desai et al. (2012) and Sevilla-Noarbe et al. (2018).
Our method uses MU_MAX -- MAG_AUTO as a proxy for SPREAD_MODEL; MU_MAX being the peak surface brightness above the background. Thus, our estimator MU_MAX -- MAG_AUTO is related to the concentration of light at the peak versus the total magnitude. This parameter has been used in Jauzac et al. (2012), Sharon et al. (2022), and Estrada et al. (2023) as input for the S/G classification.
Before flight, the point-like probability has been estimated in the plane of Fig. 9 ($x_{m}=\texttt{MAG},y_{m}=\texttt{MU\_MAX}-\tt{MAG}$) using official simulated data by merely counting the number of galaxies and stars in each bin of this 2D space. Afterwards, the point-like probability of any real observation is computed by interpolating in this 2D array at the ($x_{m},y_{m}$) position of the observation.
As shown in Fig. 9, the S/G separation is only performed in VIS-detected objects. The current rule-of-thumb to select stars in the MER catalogue is: (a) select very bright (even saturated) sources at VIS magnitudes $I_{\scriptscriptstyle\rm E}<17$; and (b) select POINT_LIKE_PROB $>0.96$. Note that the POINT_LIKE_FLAG is defined as POINT_LIKE_PROB $>0.96$ and DET_QUALITY_FLAG $=0$ so this flag can also be used. A more refined object classification, differentiating between star, galaxies and QSO, is performed by PHZ (Euclid Collaboration: Tucci et al. 2025).
Figure 9: VIS and NIR detections in the MU_MAX -- MAG_AUTO plane for real data from tile 102021495. Stars are prominently present in the bottom horizontal branch. Black crosses are sources identified as stars and the colours code the probability of being a point source.
8 The MER catalogue
The last step of the MER pipeline merges all the measurements performed in the previous steps and creating the output catalogues. The format of the MER PF output catalogues comply with the Euclid SGS standard. The main output of the cataloguing PE is the following:
  • EUC_MER_FINAL-CAT: a FITS table, storing the official Euclid source ID, i.e., the OBJECT_ID, the source coordinates, the quality flags, and all the information related to the photometry measurements. Some DEEP regions could contain imaging data from additional EXT filters. In those cases the additional EXT photometry measurements are stored in a separated FITS table, namely the EUC_MER_FINAL-DEEP-CAT.
  • EUC_MER_FINAL-MORPH-CAT: a FITS table storing all the morphological measurements performed in the MER pipeline. Each source in this table shares the OBJECT_ID information with all the other catalogues produced within this step.
  • EUC_MER_FINAL-CUTOUTS-CAT: a FITS table containing information on the corners of the source cutouts. As for the EUC_MER_FINAL-MORPHO-CAT, all the sources share the OBJECT_ID information with the rest of the output catalogues.
  • EUC_MER_FINAL-SEGMAP: a FITS binary image representing the map showing the connected pixels of the objects detected on the corresponding detection mosaics (VIS and NIR).
Appendix A gives a the detailed description of the content of the MER PF output catalogues.
External data coverage is heterogeneous across the different sky regions observed by Euclid. In particular, the Q1 release focuses on the processing of the EWS, including data from the DES (South) and UNIONS (North, Gwyn et al. in preparation) surveys, namely WHIGS, CFIS, Pan-STARRS, and WHISHES. Table 3 shows the list of bands available in the Q1 release.
Table 3: List of photometric bands available in the Q1 release. The ‘Filter label’ columns, shows how the filter is labeled in the MER catalogue data model.
Survey Band Filter label
EWS $I_{\scriptscriptstyle\rm E}$ VIS
EWS $J_{\scriptscriptstyle\rm E}$ J
EWS $H_{\scriptscriptstyle\rm E}$ H
EWS $Y_{\scriptscriptstyle\rm E}$ Y
EWS Stack of $J_{\scriptscriptstyle\rm E}$, $H_{\scriptscriptstyle\rm E}$ and $Y_{\scriptscriptstyle\rm E}$ NIR_STACK
UNIONS CFIS $u$ U_EXT_MEGACAM
UNIONS CFIS $r$ R_EXT_MEGACAM
UNIONS Pan-STARRS $i$ I_EXT_PANSTARRS
UNIONS WHIGS $g$ G_EXT_HSC
UNIONS WISHES $z$ Z_EXT_HSC
DES $g$ G_EXT_DECAM
DES $r$ R_EXT_DECAM
DES $i$ I_EXT_DECAM
DES $z$ Z_EXT_DECAM
8.1 E(B-V) estimates
The distribution of galaxies on large scales probed by the Euclid satellite is affected by Galactic absorption and therefore the study of the galaxy clustering has to be corrected by the different survey depths, which could induce artefacts in the analysis. Moreover, the errors in this correction should be taken into account in the covariance matrix.
OU-MER provides the extinction $E(B-V)$ as estimated by the R1.20 Planck data release and estimates an error in the $E(B-V)$. This is done via the LE3_GALEXT_ED PF, a Python program that imports the public healpy [^6] https://healpix.jpl.nasa.gov/ in order to read the extinction map described in Planck Collaboration (2014). The map provides values for $E(B-V)$ and $\tau$, the optical depth of dust as measured at 353 GHz, together with its associated error $\tau_{\text{err}}$. The error associated with $E(B-V)$ can be calculated via the ratio between $\tau_{\text{err}}$ and $\tau$,
$E(B-V)_{\text{err}}=E(B-V)\,\frac{\tau_{\text{err}}}{\tau}.$ (18)
The algorithm takes a list of sky positions in a specified coordinate system and computes the dust extinction $E(B-V)$ based on the reference extinction map, assigning an $E(B-V)$ measurement and its associated error to each source in the catalogue.
8.2 Identification of spurious sources
[…]
  • the importance ranking of each input feature, which determines the most relevant properties in classifying the sources as spurious or real, resulting in the ground bands $g$, $r$, $i$, and $z$ obtaining the highest ranking;
  • the actual spurious probability (SPURIOUS_PROB) for each object in the photometric catalogue.
Given SPURIOUS_PROB, we identify a source to be spurious by adopting a probablity threshold of $50\%$. With such a threshold, we identify up to $90\%$ of the true spurious sources in our test sample, while only $0.13\%$ of real galaxies or stars are wrongly labelled as spurious. A more conservative threshold of $75\%$ allows us to increase the purity (only $0.04\%$ of real sources are lost), at the cost of a lower completeness of the spurious sample (only $80\%$ of them are correctly identified and removed). We found $50\%$ to be a good compromise between purity and completeness, and this represents our default choice.
8.3 Bright star masking
Figure 10: The bright star masks (regions marked in red) in two small cutouts of the $I_{\scriptscriptstyle\rm E}$ (left) and $H_{\scriptscriptstyle\rm E}$ (right) band image of tile 102021017. The central ring and the length of the diffraction and blooming features scale with brightness of the star.
As a survey satellite, Euclid observes bright stars up to 8th magnitude brightness. The image of bright stars is characterised by a large and often saturated centre, together with diffraction spikes and possibly a blooming signature from the VIS detectors. These features are difficult to handle in the detection, deblending, and photometry. To mitigate this we map the contours of bright stars on the co-added VIS and NIR images as polygons.
Starting from positions provided by Gaia we map the imprint of stars on the individual VIS and NIR calibrated images using simple geometric forms:
  • a central circle for the stellar disk;
  • rectangles for the three diffraction spikes;
  • a rectangle for the VIS blooming pattern.
The size of these simple forms scales with the brightness of the stars. For every Gaia star we build the union of the simple forms as a polygon and then project the polygon from each calibrated image to the co-added mosaic. The final polygon for each Gaia star is then set as the union of the polygons projected from the individual VIS and NIR calibrated images. As output we provide for all co-added measurement images a list of the combined polygons for the Gaia stars on that image as a JSON [^7] https://www.json.org/json-en.html file.
Figure 10 shows with the red lines the bright star masks in two small cutouts of the $I_{\scriptscriptstyle\rm E}$ (left) and $H_{\scriptscriptstyle\rm E}$ (right) band image of tile 102021017. The different colour denote the brightness bracket of star, which scales the size of the geometric forms constituting the masks. All objects within the bright star masks in $I_{\scriptscriptstyle\rm E}$ or any of the bright star masks in $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$ or $H_{\scriptscriptstyle\rm E}$ are marked in the column DET_QUALITY_FLAG in two separate bits.
8.4 Matching of Gaia sources
We automatically match the Euclid objects with the Gaia DR3 sources (Gaia Collaboration et al. 2023). To optimise the matching result, we propagate the positions of the Gaia objects from the DR3 epoch J2016.0 to the reference time of the MER catalogue, which is averaged from the VIS and NIR detection image reference times. The maximum allowed distance of $0\overset{\prime\prime}{.}3$, and the ID of the matched Gaia source is stored in the catalogue column GAIA_ID. The column GAIA_MATCH_QUALITY stores the matching quality as the squared distance between the Gaia and Euclid positions. Depending on Galactic latitude the number of Gaia matches is between $800$ and $3000$ objects in the final catalogues.
9 Validation pipeline
The ultimate purpose of the validation is to assure that output data of the MER pipeline are within the specification and requirements and contribute to reach the science goals of the Euclid project. Data validation helps also to identify processing issues or in general data quality problems. In the OU-MER validation we measure key properties of the MER outputs, such as object properties and image characteristics. If possible we compare and extrapolate the object properties with ground-truth information provided by the Gaia satellite.
The OU-MER validation is run automatically as a separate pipeline directly after the processing pipeline. The Q1 dataset is made up of three distinct fields, namely EDF-N, EDF-S, and EDF-F (Euclid Collaboration: Aussel et al. 2025), covered by several tiles each. We aggregate the information of all the tiles covering a given field, in order to analyse the statistical trend of the parameters computed in the validation pipeline. This aggregated information is an essential tool to monitor and improve the MER processing.
In the following Sections we present the most important validation results either for individual selected tiles or the aggregate numbers for a dataset such as EDF-S.
9.1 Astrometry
To assess the accuracy of the MER pipeline astrometry, we measure, for each tile, the mean positional offset with respect to Gaia (Fig. 11). Tiles with larger values of FILLING_FACTOR (Sect. 9.3), with a comparably large offset are located at the border of the field. This is expected since both these offsets originate from systematic residuals in the VIS astrometry (Euclid Collaboration: McCracken et al. 2025), which are not evenly sampled at the border.
Figure 11: Distribution of the astrometric offsets for all tiles in the EDF-S field.
9.2 Image Depth
Figure 12: The $5\,\sigma$ depth depth measured in all imaging data of tile 102021017 using flux measurement in $2^{\prime\prime}$ apertures in the sky background.
We determine the depth of the imaging data by measuring the flux in several thousand randomly selected circular apertures. We restrict the measurements to background areas by skipping over apertures that have non-zero values in the segmentation image or contain empty image areas. From the scatter of the flux values in these apertures we compute the image depth in the corresponding filter.
As an example Fig. 12 shows the $5\,\sigma$ depth in the multi-band data of tile 102021017 measured in apertures of $2^{\prime\prime}$ diameter. We have chosen a $2^{\prime\prime}$ diameter since corresponds to the typical $2\,{\rm FHWM}$ diameter in the worst resolution band (see Sect. 6). The depth estimates in other Q1 tiles are very similar to the ones shown in Fig. 12, especially for the data observed with the Euclid satellite $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$.
Note that these depth estimates are done on aligned images that are resampled from their original pixel scale, which certainly has an effect on these measurements. Also the variable exposure time within fully covered images and especially in the border tiles of Q1 may not fully be represented in a single depth value per band. Consequently we do not see these estimates as very accurate and absolute values. We use them as relative estimates for a set of tiles such as that released in Q1 to control the data quality.
9.3 Detection
In this part we collect for each tile all characteristic numbers associated to the object detection, cross-matching and object marking. We compute the object density in the HealPix pixels of the tile core area and determine the FILLING_FACTOR of the tile, which is the fraction of HealPix pixels containing at least one detected object. For interactive checks we illustrate these numbers in images such as Fig. 13, which shows the positions of the VIS and NIR detect numbers for tile 102021017 in EDF-S on the left and right panel, respectively. The bright star masked objects as well as the objects identified from Gaia are marked in both panels. The vertical “lines” of NIR detected objects in the right panel are caused by an insufficient masking of pixels affected by persistence (see Euclid Collaboration: Polenta et al. 2025). For an entire Q1 dataset the detection validation numbers help to identify border tiles with a small FILLING_FACTOR or data quality issues such as shallow or incomplete coverage in a detection band when comparing the number of VIS and NIR detected objects.
Figure 13: The positions of all detected objects in tile 102021017. The color mark the various object categories as indicated on the right.
9.4 PSF
The PSF propagation method described in Sect. 5 is validated using Gaia stars with CLASSPROB_DSC_COMBMOD_STAR ¿ 0.99. In addition, a magnitude cut of MAG_G_GAIA ¿ 18 is applied to avoid possible saturation effects (see Gaia DR3 Data Model for details on the Gaia source catalogue definitions). For Q1, depending on the field, this results on a typical sample of 300 to 600 Gaia stars per MER tile.
We extract stamps in the MER mosaics around the Gaia star positions corrected for their proper motion. We use small stamps, $2\overset{\prime\prime}{.}1\times 2\overset{\prime\prime}{.}1$ for the Euclid bands and $4\overset{\prime\prime}{.}5\times 4\overset{\prime\prime}{.}5$ for the EXT mosaics, to minimise contamination from nearby sources. The star stamps are then fitted to a two-dimension Gaussian function and the fit FWHM values are compared with the propagated PSF stamps at the star positions. PSF stamps are fitted with exactly the same method. Ideally one would expect a 1 to 1 correlation between the two fitted values. Any systematic shift or large scatter in the correlation would imply that either the propagation method is not correct or that the input PSFs are not applicable to the input image data.
Figure 14 shows the propagated VIS PSF FWHM values for an example tile in the EDF-S field. The FWHM could change by more than 20%, depending on the source spatial position. Large-scale variations can be attributed to PSF changes in the VIS instrument field of view (Euclid Collaboration: McCracken et al. 2025). Figure 15 shows the comparison between the propagated PSF stamps and the Gaia stars extracted from the VIS mosaic. The agreement is very good with minimal excursions from the the ±$\pm 2\%$ offset lines. As discussed in Sect. 5, the results for the NIR bands are not as good, but they should improve in DR1 with newer NIR input PSFs. The agreement for the EXT bands is generally also very good. In these cases the PSF spatial variations can be attributed to varying seeing conditions in the images that were used to produce the mosaics.
Figure 14: Spatial variation of the MER catalogue-PSF for the VIS band on an example MER tile in the EDF-S field (tile index 102021019). FWHM values are measured fitting a two-dimensional Gaussian model to each propagated PSF stamp. Large-scale variations can be traced back to PSF changes along the VIS focal plane (see Euclid Collaboration: McCracken et al. 2025, for more details). In this example, the sharp FWHM change at RA $\sim$ 65°65°65  ° coincides with the transition between two different Euclid observation.
Figure 15: Comparison between the Gaia star profiles extracted from the VIS mosaic in tile 102021019 and the propagated VIS PSFs stamps at the Gaia star positions. FWHM values are measured fitting a two-dimensional Gaussian model to the PSF stamps and the Gaia star profiles. The identity line (continuous) and ±$\pm 2\%$ FWHM difference lines (dashed) show that both measurements agree within 2% in most of the cases.
9.5 Photometry
The photometric accuracy of the MER pipeline in each of the three Q1 datasets can be assessed by means of the magnitude offsets provided by the MER Gaia validation. The offsets are computed for each band (namely $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ along with the EXT-UNIONS filters for the case of EDF-N, and $I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$, and $H_{\scriptscriptstyle\rm E}$ plus the EXT-DECAM filters for the case of EDF-N and EDF-F) and for two different photometric approaches (APHOT and TPHOT), taking into account the non-saturated sources with a magnitude $I_{\scriptscriptstyle\rm E}<19.5$ that have been detected by the MER pipeline and that have been successfully cross-matched against the Gaia DR3 catalogue.
A photometric offset can be calculated for each source using a set of transformation functions generated to provide an expected analytic value for the magnitude difference between a given Euclid band and the total $g$ magnitude in the Gaia system. Finally, the median magnitude offset of the all selected sources is computed, in order to have an aggregate quantity to describe the overall photometric accuracy per tile. For each of the three Q1 fields, this mean magnitude offset was computed after applying a $3\,\sigma$-clipping over the per-tile values, a procedure that also allowed the identification of outlier tiles where the specific magnitude offset value was not aligned with the global trend of the field (this typically being a direct consequence of major distorting elements that affected the normal execution of the MER pipeline). We find overall reasonable agreement, with offsets at the percent level varying from band to band. For instance, as it can be observed in Fig. 16 for the case of EDF-S, we find mean magnitude offsets of $-0.0688\pm 0.0029$ mag in $I_{\scriptscriptstyle\rm E}$, $-0.0202\pm 0.0100$ mag in $Y_{\scriptscriptstyle\rm E}$, $-0.0306\pm 0.0109$ mag in $J_{\scriptscriptstyle\rm E}$ and $-0.0078\pm 0.0128$ mag in $H_{\scriptscriptstyle\rm E}$, respectively in what concerns aperture photometry (APHOT), while, for the case of template fitting photometry (TPHOT), we report mean magnitude offsets of $-0.0809\pm 0.0128$ mag in $Y_{\scriptscriptstyle\rm E}$, $-0.0746\pm 0.0131$ mag in $J_{\scriptscriptstyle\rm E}$ and $-0.0426\pm 0.0124$ mag in $H_{\scriptscriptstyle\rm E}$, respectively. For the case of the EXT-DECAM bands included in EDF-S, we find mean magnitude offsets of $-0.0285\pm 0.0142$ mag in $g$, $-0.0086\pm 0.0054$ mag in $r$, $0.0007\pm 0.0072$ mag in $i$ and $-0.0018\pm 0.0102$ mag in $z$, respectively in what concerns APHOT, while for the case of TPHOT, we report mean magnitude offsets of $-0.0103\pm 0.0208$ mag in $g$, $0.0188\pm 0.0059$ mag in $r$, $0.0275\pm 0.0087$ mag in $i$ and $0.0246\pm 0.0126$ mag in $z$, respectively. However, we point out that these offsets values have been continuously changing in the various phases of the observational campaign. Their causes are under investigation, with difference in photometric techniques, PSF, and calibration related issues being the most probable culprits. More detailed information, including data for other fields and bands, can be found in the Q1 release notes associated to this paper (see Sect 9.7).
It should be kept in mind that all of the validation involves bright stars only, which are not the main targets of the survey and therefore of the pipeline; checks against archival catalogues of galaxies have been performed, yielding reasonable agreement, but the uncertainties are too large to allow for a thorough evaluation of the accuracy. We have verified with random direct checks on the final mosaics that while for bright stars the aperture photometry estimate yields the best estimates, template fitting performs better for extended objects, not only helping to beat blending issues, but providing more accurate measurements in all cases for galaxies.
Figure 16: Mean magnitude offsets for aperture (A-PHOT) and template fitting (T-PHOT) photometry in EDF-S derived from the results computed by the validation pipeline against cross-matched Gaia sources for the Euclid ($I_{\scriptscriptstyle\rm E}$, $Y_{\scriptscriptstyle\rm E}$, $J_{\scriptscriptstyle\rm E}$ and $H_{\scriptscriptstyle\rm E}$) and EXT-DECAM bands ($g$, $r$, $i$ and $z$). The error bars correspond to the specific standard deviation associated to each mean magnitude offset calculated.
9.6 Point-like probability
In order to identify possible errors in the identification of point-like objects, we compare for each tile the number of Gaia matches with the number of objects with POINT_LIKE_PROB $>0.96$. As can be seen in Fig. 17, these two parameters are roughly equal. In EDF-N (the galactic latitude is $030\overset{\circ}{.}$ 0), they can be both small if the tile is not fully observed, or both around 2000 if the filling factor is close to 1. Any strong deviation from the dashed gray line ($y=x$) indicates a tile that needs further analysis (it is often a good indicator that some solar X ray burst has hit the VIS detectors).
[…]
At the tile level we validate POINT_LIKE_PROB against Gaia matches identified in the MER final catalog (See Sect. 17) as shown in Fig. 18 for tile 102021017. There are 1104 objects with Gaia matches and 1070 with a high point-like probability (POINT_LIKE_PROB $>0.96$ ). From these 1070 there are 744 with and 326 without Gaia matches. The not matched objects are concentrated beyond the limiting magnitude of Gaia.
[…]
9.7 Q1 release notes and known issues
Q1 data come with a list of caveats and known issues identified also thanks to the effort of EC scientists outside the perimeter of OU-MER and, in general, the SGS. An exhaustive description of how to deal with OU-MER products in the Q1 dataset can be found in the Q1 Explanatory Supplement [^8] https://euclid.esac.esa.int/dr/q1/expsup.
10 Conclusions and outlook
In this paper we have presented a general overview of the OU-MER pipeline, with a particular focus on the work done for the data released as part of Q1. The Q1 exercise has allowed us to identify some areas in which to focus our work towards future, larger data releases such as DR1.
The main goal of OU-MER is to provide photometric and morphological measurements to both the customer OUs in the Euclid SGS (OU-PHZ, OU-SIR, OU-SHE) and the general scientific community. The main effort towards the upcoming Data Release 1 (DR1) is to improve the consistency of our photometry. The validation pipeline has shown that the current results for photometry are reasonably good, although with some room for improvement. As part of the continuing pipeline development we are currently conducting various tests to reach a full understanding of some issues that are outlined in the Q1 release notes (details in the Q1 Explanatory Supplement linked in Sect. 9.7).
The most important assessment on the quality of the OU-MER pipeline output products is going to come from the downstream units such as OU-PHZ, -SHE and -LE3 when using larger quantities of Euclid data in a cosmology analysis such as planned for DR1 and later releases.
Acknowledgements.
The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org).
This work has made use of the Euclid Quick Release (Q1) data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b.
Based on data from UNIONS, a scientific collaboration using three Hawaii-based telescopes: CFHT, Pan-STARRS, and Subaru www.skysurvey.cc .” Correspondingly, please add “Based on data from the Dark Energy Camera (DECam) on the Blanco 4-m Telescope at CTIO in Chile https://www.darkenergysurvey.org .” if you use MER data from the southern hemisphere. If you use Gaia data, please acknowledge: “This work uses results from the ESA mission Gaia, whose data are being processed by the Gaia Data Processing and Analysis Consortium https://www.cosmos.esa.int/gaia
Appendix A The OU MER Data Model
All Euclid data products must comply with the SGS Data Model (DM) and are documented in the MER Data Product Description Document [^9] https://euclid.esac.esa.int/dr/q1/dpdd (DPDD). The DPDD also states, for each product, if the product itself is ingested into the Euclid Science Archive System (SAS, Euclid Collaboration: Aussel et al. 2025). In this section we provide a brief summary of the OU-MER pipeline deliverables.
A.1 Output images and segmentation maps
Together with the photometric and morphological catalogues, the OU-MER pipeline outputs the set of images on which the photometric and morphological analysis is performed. Those images areas follows.
  • The background-subtracted mosaics, one for each input band. The associated data product in the DPDD is DpdMerBksMosaic.
  • The mosaics used to perform the object detection, one for VIS (VIS) and one for the stack of the three NIR bands (NIR_STACK). The associated data product in the DPDD is DpdMerDetectionMosaic.
  • A map showing the connected pixels of the objects detected on the corresponding detection mosaics (VIS+NIR_STACK). The associated data product in the DPDD is DpdMerSegmentationMap.
Note that the segmentation map, together with the catalogue, refers to the inner region of a MER tile (see Sect. 3.1), while the mosaics cover the whole tile region. Figure 19 shows the inner and outern tile region of a MER tile with respect to a reference mosaic.
Figure 19: Outer and inner tile regionsrelative to an OU-MER mosaic. The green frame shows the MER tile outern region, while the red frame is the inner MER tile region. The segmentation map and the catalogue refer to the inner tile region; all the sources between the outer and inner regions are not catalogued. A value of $-1$ is assigned to pixels related to sources outside the inner tile region. A standard MER tile (outern region) covers an area of $32^{\prime}\times 32^{\prime}$. The border between the outern and inner region has a depth of $2^{\prime}$.
A.2 Naming convention and data format
The names of the data files associated to the MER data products are formed using the following format:
EUC_MER_[TYPE]_TILE[INDEX]-[RND]_[TIME].[EXT]
where TYPE is one of the file types described below, INDEX is the associated MER tile index (e.g., 102022523), RND is a random string that it is used to avoid possible file name duplications (e.g., 15E4F0), TIME is the file creation timestamp (e.g., 20241121T114727.022303Z) and EXT is the file extension (fits.gz or json).
The following are the file types used in the MER PF (see Table 3 for a complete list of band label values):
  • BGSUB-MOSAIC-<band>: Mosaic image data.
  • MOSAIC-<band>-RMS: Mosaic rms data.
  • MOSAIC-<band>-FLAG: Mosaic flag data.
  • BGMOD-<band>: Mosaic subtracted background model.
  • FILTER-TRANSMISSION-<band>: Mosaic average filter transmission wavelength.
  • DETECTOR-LAYERING-<band>: Mosaic input layers information.
  • CATALOG-PSF-<band>: Mosaic catalogue PSF.
  • GRID-PSF-<band>: Mosaic grid PSF.
  • STAR-MASKS-<band>: Mosaic bright star spatial masks.
  • FINAL-SEGMAP: Detection segmentation map.
  • FINAL-CAT: Catalogue with the main photometry information.
  • FINAL-MORPH-CAT: Catalogue with the morphology information.
  • FINAL-CUTOUTS-CAT: Catalogue with the source cutout coordinates.
  • FINAL-DEEP-PHOTO: Catalogue with extra photometry information available only on the Euclid DEEP fields.
The DPDD provides more details about the content of the different MER data files.
A.3 Column-by-column catalogue description
The catalogues produced by OU MER pipeline list more that 3 20032003\,2003 200 columns. The structure of the photometry catalogues (EUC_MER_FINAL-CAT and EUC_MER_FINAL-DEEP-CAT) is summarised in sections Main catalogue FITS file and DEEP field photometry catalogue FITS file of the MER DPDD. Section Morphology catalogue FITS file describes the morphology catalogue (EUC_MER_FINAL-MORPHO-CAT), while section Cutouts catalogue FITS file lists the columns available in the cutout catalogue (EUC_MER_FINAL-CUTOUTS-CAT).